# FML-bench: A Benchmark for Automatic ML Research Agents Highlighting the Importance of Exploration Breadth

### v1

@@ 4.2 Unified Input-output Interface
A core design of our benchmark is to support direct utilization of machine learning research GitHub repositories. Real-world repositories vary significantly in execution pipelines, output formats, and evaluation protocols. We solve this through unified input-output interfaces that preserve repository complexity.
@@
Repository outputs are various (e.g., from text files to JSON formats). However, most outputs share a common structure: performance metrics on specific datasets. We provide a post-processing module that converts diverse task outputs into a standardized format, enabling consistent metric extraction across all tasks while preserving native output mechanisms.
This design bridges the gap between evaluation standardization and real-world repository diversity, enabling rigorous assessment of agents on practical code optimization tasks.
@@ 5.1 Settings
Benchmark Task Settings
Our benchmark encompasses 8 fundamental machine learning challenges, each implemented using established repositories with baselines: 1) Generalization using the DomainBed (Gulrajani & Lopez-Paz, 2020) repository with ERM (Vapnik, 1998) baseline, optimizing out-of-domain accuracy; 2) Data Efficiency using the Easy-Few-Shot-Learning (Sicara, 2024) repository with Prototypical Networks (Snell et al., 2017) baseline, maximizing few-shot classification accuracy; 3) Representation Learning using the Lightly (Lightly-AI, 2025) repository with MoCo (He et al., 2020) baseline, maximizing linear probing accuracy on frozen encoders; 4) Continual Learning using the Continual-Learning (van de Ven et al., 2022) repository with Synaptic Intelligence (Zenke et al., 2017) baseline, optimizing average accuracy across sequential tasks; 5) Causality using the CausalML (Chen et al., 2020) repository with DragonNet (Shi et al., 2019) baseline, minimizing absolute error in average treatment effect estimation; 6) Robustness and Reliability using the Adversarial Robustness Toolbox (ART) (Nicolae et al., 2018) repository with dp-instahide (Borgnia et al., 2021) defense baseline, optimizing defense scores against backdoor attacks; 7) Privacy using the PrivacyMeter (Murakonda & Shokri, 2020) repository with Wide-ResNet-28-2 (Zagoruyko & Komodakis, 2016) baseline, minimizing membership inference attack AUC; and 8) Fairness and Bias using the AIF360 (Bellamy et al., 2018) repository with Adversarial Debiasing baseline, minimizing absolute average odds difference while preserving classification performance. Comprehensive details are provided in Appendix B.
@@
Each agent is required to execute in three independent rounds. In each round, the agent is assigned a fixed budget of total steps = 100 (iterations). We select the best result achieved among the three rounds based on the target metric computed on the test set.
Evaluation Metrics
We evaluate agent performance on our benchmark using the proposed metrics: Utility, Diversity, Academic Contribution Rate, Cost, and Step Success Rate. In addition, we report the Step Completion Rate, which calculates the proportion of executed steps relative to the required total, since AIDE and Claude Code may exhibit premature termination. See Sec. C in the appendix for detailed calculation of metrics. We present detailed Utility results across the 8 tasks (Tab. 2), and report averages of the other metrics over the 8 tasks (Tab. 3); full per-task results are provided in Appendix E.
@@ 5.2 Implementation Details
To enable these agents to operate on our benchmark, several modifications were necessary. We adapted TheAIScientist by fixing compatibility issues and extending its functionality to support the requirements of our benchmark, such as executing experiments in real repositories and reporting results consistently. As for AIDE, we employed its cloud-based commercial variant, Weco, as the operational interface, adapting our benchmark to integrate with its workflow despite limited control over its internal mechanisms. For Claude Code, we designed a prompting scheme (see Appendix F) that enabled it to function as an automatic research agent, capable of reading code, generating hypotheses, and proposing modifications grounded in experimental feedback.
@@
We employ GPT-5 (2025-08-07) and Gemini-2.5-Pro (2025-06-17) for TheAIScientist and AIDE. For Claude Code, it is constrained to its native models, and therefore, we use Opus-4.1 (2025-08-05).
@@ 5.3 Results and Discoveries
5.3.1 Comparison of Agents with Different Research Strategies
As shown in Tab. 2, the combination of TheAIScientist with Gemini-2.5-Pro achieved the best performance, ranking first in 4 out of 8 tasks. The combination of TheAIScientist with GPT-5 ranked second, securing top results in 2 out of the 8 tasks. These findings suggest that TheAIScientist performs better in discovering novel and effective machine learning methods, compared to AIDE and Claude Code.
As illustrated in Fig. 2, TheAIScientist adopts a research exploration strategy that is broad but shallow, while AIDE maintains both medium breadth and depth. In contrast, Claude Code exhibits a narrow yet deep exploration pattern. For a detailed explanation of the research strategies adopted by each agent, refer to Section 5.1. Considering the research exploration strategy together, the results suggest that a broader research exploration space adopted by TheAIScientist is more effective for discovering promising ideas. This insight offers practical guidance for real-world research: broadly exploring diverse ideas could be more productive than focusing on a single direction.
Table 2: Comparison of best performance (Utility) among different agents. G2.5-Pro denotes Gemini-2.5-Pro. $(\uparrow)$ indicates higher is better, $(\downarrow)$ indicates lower is better.
@@ 5.3.1 Comparison of Agents with Different Research Strategies
Causality $(\downarrow)$ 0.2057 0.0434 0.0058 0.0155 0.0000 0.0063
Robust. & Rel.()($\uparrow$) 0.6966 0.9623 0.9570 0.9708 0.9653 0.9469
@@
Fair. & Bias $(\downarrow)$ 0.1471 0.0002 0.0053 0.0004 0.0030 0.0149
Table 3: Comparisons of different agents across diversity, academic contribution rate, cost, and success metrics. Contrib. denotes Contribution and Cons. stands for Consumption.
@@
Diversity $(\uparrow)$ 24.95±9.63 20.66±10.85 20.26±9.37 18.49±14.87 12.02±8.04
Academic Contrib. Rate $(\uparrow)$ 0.83±0.27 0.78±0.24 0.84±0.20 0.65±0.34 0.25±0.36
Total Token Cons. (M) $(\downarrow)$ 6.07±2.05 5.33±2.55 0.85±0.59 2.59±0.90 8.32±3.10
Step Token Cons. (M) $(\downarrow)$ 0.06±0.02 0.05±0.03 0.03±0.03 0.04±0.03 1.77±1.32
Total Time Cons. (H) $(\downarrow)$ 8.32±3.15 12.23±7.93 2.46±1.77 8.06±6.44 1.27±1.77
Step Time Cons. (Min) $(\downarrow)$ 10.39±10.79 12.73±14.64 10.79±12.86 8.65±7.92 12.49±10.77
Step Success Rate $(\uparrow)$ 0.92±0.08 0.87±0.12 0.65±0.25 0.73±0.22 0.83±0.18
Step Completion Rate $(\uparrow)$ 1.00±0.00 1.00±0.00 0.38±0.22 0.79±0.25 0.07±0.05
5.3.2 Analysis of Diversity
As shown in Tab. 3, TheAIScientist shows the highest average diversity (GPT-5: $24.95\pm 9.63$; Gemini-2.5-Pro: $20.66\pm 10.85$). AIDE is lower but comparable (GPT-5: $20.26\pm 9.37$; Gemini-2.5-Pro: $18.49\pm 14.87$), while Claude Code is markedly lower ($12.02\pm 8.04$). This pattern mirrors how the agents explore solutions. TheAIScientist advances several ideas in parallel. AIDE grows ideas through a tree of iterative refinements. Claude Code, by contrast, tends to proceed along a single linear track. By comparison, parallel exploration broadens the search and produces higher measured diversity, while linear iteration confines the search and curbs diversity.
We further analyze the relationship between code diversity and task performance using the correlation coefficient $r$ (details are shown in Appendix Fig. 4). The results suggest that, overall, diversity is positively correlated with performance. Specifically, among the 8 tasks, 4 tasks show strong positive correlations, 2 weak positive correlations, and 2 negative correlations. The most notable effects appear in Continual Learning ($r=0.96$), Fairness & Bias ($r=0.86$), and Generalization ($r=0.72$), with a moderate correlation also observed in Data Efficiency ($r=0.48$). These findings indicate that higher code diversity tends to be associated with improved task performance, although the strength of this relationship varies by task.
5.3.3 Academic Contribution Rate
The academic contribution rate provides further insights into the characteristics of each agent. It helps disentangle the impact of academic value from other factors such as engineering effort and diversity. According to Tab. 3, TheAIScientist generally exhibits a slightly higher academic contribution rate than AIDE, whereas Claude Code consistently shows the lowest rate. This suggests that the ideas and code modifications proposed by TheAIScientist are more closely aligned with methodological advancements, rather than relying on engineering tricks to boost performance. Furthermore, comparing GPT-5 and Gemini-2.5-Pro reveals that Gemini-2.5-Pro tends to propose more engineering-oriented solutions than GPT-5. In summary, for automatic ML research agents, we prioritize agents that can generate hypotheses with strong academic value while simultaneously delivering better Utility $U$ (higher performance).
@@ 5.3.5 Additional Observations
Characteristics of Claude Code
Since all actions are executed based on LLM decisions rather than fixed procedures, Claude Code often fails to follow prompt instructions, frequently terminating experiments prematurely. Despite this, it demonstrates a high improvement speed (refer to the improvement curves Fig. 3 and speed comparison Tab. 12 in the Appendix). In addition, its academic contribution rate is low, with a strong emphasis on engineering. This may be attributed to its nature as a general-purpose agent, rather than a specialized automatic ML research agent.
@@
We found that AIDE sometimes misinterprets the structure and logic of target codebases. In certain cases, it generated new classes or components that were never integrated into the actual execution pipeline, resulting in no functional improvement over the baseline. As shown in Tab. 2, AIDE failed to improve the baseline in tasks related to Generalization and Data Efficiency. This may stem from the fact that AIDE only supports iterative modifications on a single file. However, real-world ML research codebases are often complex and span multiple files, making AIDE insufficient for addressing realistic research tasks.
@@
We encountered the issue of early termination of AIDE and Claude Code. For AIDE, the agent sometimes terminated prematurely due to its commercial version, Weco, which relies on cloud infrastructure that occasionally failed during execution. For Claude Code, early stopping was often triggered by the model’s internal reasoning, where the LLM would decide not to continue even when further actions are possible.
@@ 6 Conclusion
In this work, we introduce FML-bench, a benchmark that assesses automatic machine learning research agents on 8 diverse and fundamental ML problems drawn from real-world codebases and authentic research workflows. Within the benchmark, we introduce a five-dimensional evaluation protocol that facilitates a more systematic assessment, moving beyond exclusive reliance on final performance. Building on FML-bench, we conducted a systematic analysis of three state-of-the-art automatic research agents. Our findings reveal that agents capable of generating and evaluating multiple hypotheses across diverse directions, such as TheAIScientist, tend to outperform those that focus on iteratively refining a single line of thought (e.g., Claude Code). These findings suggest that a broader exploratory capacity contributes more significantly to overall success once a basic level of proficiency is established. Overall, FML-bench provides a robust and practical foundation for evaluating the capabilities of research agents and offers a pathway toward building more effective, generalizable, and scientifically productive research agents.
Reproducibility Statement
The paper describes the benchmark design, evaluation protocol, experimental settings, and implementation details, including baseline configurations and prompt specifications. All benchmark code, experiment prompts, and configuration files are open-sourced and available at: https://github.com/qrzou/FML-bench.
@@ Acknowledgements
Appendix
Appendix A Use of LLMs
In this study, we employed LLMs (ChatGPT and Claude AI) to assist with manuscript writing. Specifically, these tools were used to polish the writing (e.g., improving fluency, grammar, and clarity of expression), aid in formatting, and support proofreading. Additionally, we used them to facilitate the retrieval of related work by suggesting potentially relevant literature.
Appendix B Task Settings
Generalization
@@ Appendix B Task Settings
We adopt the DomainBed (Gulrajani & Lopez-Paz, 2020) repository with Empirical Risk Minimization (ERM) (Vapnik, 1998) as the baseline and evaluate on the ColorMNIST (Arjovsky et al., 2019) dataset. The evaluation metric is accuracy on a held-out domain. The agent is required to train on a source domain and generalize to a target domain under distribution shift. The task objective is to improve out-of-domain generalization performance. This task evaluates the agent’s ability to develop algorithms that transfer effectively across domains, which is critical for robust real-world deployment.
Causality
We utilize the CausalML (Chen et al., 2020) repository with DragonNet (Shi et al., 2019) as the baseline model, evaluated on a synthetic dataset constructed from Setup A in Nie & Wager (2021), Quasi-Oracle Estimation of Heterogeneous Treatment Effects. The performance metric is the absolute error in average treatment effect (ATE) estimation. The agent’s goal is to develop improved causal inference strategies that minimize this error. This task assesses the agent’s capacity to reason about interventions, which is essential for decision-making in high-stakes environments.
Data Efficiency
@@
For data-efficient learning, we employ the Easy-Few-Shot-Learning (Sicara, 2024) repository with Prototypical Networks (Snell et al., 2017) as the baseline, evaluated on the Mini-ImageNet (Vinyals et al., 2016) dataset. The agent must operate under a frozen backbone and propose improved algorithms for metric-based classification in the embedding space. Accuracy is used as the evaluation metric. This task measures the agent’s ability to enhance few-shot learning performance by optimizing distance-based reasoning under tight data constraints.
Robustness and Reliability
We adopt the Adversarial Robustness Toolbox (ART) (Nicolae et al., 2018) repository, using dp-instahide (Borgnia et al., 2021) as the defense baseline on a poisoned MNIST (Deng, 2012) dataset constructed with the “pattern” backdoor method. The evaluation metric is a defense score, defined as the harmonic mean of clean accuracy and resistance accuracy against backdoor attacks. The agent is tasked with proposing defenses that reduce the effectiveness of poisoning attacks while maintaining high clean performance. This task probes the agent’s ability to improve model robustness against adversarial corruption.
Privacy
This task uses the PrivacyMeter (Murakonda & Shokri, 2020) repository and Wide-ResNet-28-2 (Zagoruyko & Komodakis, 2016) as the baseline model trained on CIFAR-10 (Krizhevsky et al., 2009). The evaluation targets membership inference attacks, with performance measured by the area under the ROC curve (AUC). The agent must design improved defense mechanisms that minimize the AUC, thereby reducing information leakage. This task is critical for assessing the agent’s ability to enhance privacy protections in the face of inference attacks.
Representation Learning
For self-supervised learning, we employ the Lightly (Lightly-AI, 2025) repository with MoCo (He et al., 2020) as the baseline and evaluate on CIFAR-10 (Krizhevsky et al., 2009). The primary evaluation metric is the accuracy of a linear classifier trained on top of the frozen encoder after pretraining. The agent is expected to devise improved representation learning methods that yield higher linear probing accuracy. This task tests the agent’s ability to learn generalizable and semantically meaningful features from unlabeled data.
Continual Learning
@@
We use the Continual-Learning (van de Ven et al., 2022) repository with Synaptic Intelligence (SI) (Zenke et al., 2017) as the baseline, evaluated on splitMNIST (Deng, 2012) in the class-incremental learning (Class-IL) scenario. The model must learn sequentially across multiple contexts using a shared 10-way classification head. The evaluation metric is the average accuracy across all tasks. The agent is expected to develop algorithms that mitigate catastrophic forgetting and maintain performance across contexts. This task evaluates long-term adaptability in non-stationary environments.
Fairness and Bias

### v2

@@ 4.2 Unified Input-output Interface
In real-world scientific research, different ML projects often employ distinct codebases with varying execution pipelines, output formats, and evaluation protocols, making it challenging to construct a unified benchmark across diverse research problems. A core design of FML-bench addresses this challenge through unified input-output interfaces that accommodate repository diversity while preserving their inherent complexity. This design enables our benchmark to support diverse ML research tasks built upon existing repositories, bringing evaluation closer to real-world scientific research settings.
@@
Repository outputs are various (e.g. from text files to JSON formats). However, most outputs share a common structure: performance metrics on specific datasets. We provide a post-processing module that converts diverse task outputs into a standardized format, enabling consistent metric extraction across all tasks while preserving native output mechanisms. This design bridges the gap between evaluation standardization and real-world repository diversity, enabling rigorous assessment of agents on practical code optimization tasks.
Evaluation Integrity
Our benchmark implements protective measures to ensure evaluation integrity. Specifically, we protect evaluation files as read-only and prevent agent modification. Agents must utilize these protected evaluation files to assess their proposed methods.
@@ 5.1 Settings
Each agent is required to execute in three independent rounds. In each round, the agent is assigned a fixed budget of total steps = 100. We select the best result achieved among the three rounds based on the target metric computed on the test set.
@@ 5.2 Implementation Details
To enable these agents to operate on our benchmark, several modifications were necessary. We adapted TheAIScientist by fixing compatibility issues and extending its functionality to support the requirements of our benchmark, such as executing experiments in real repositories and reporting results consistently. As for AIDE, we employed its cloud-based commercial variant, Weco, as the operational interface, adapting our benchmark to integrate with its workflow despite limited control over its internal mechanisms. For Claude Code, we designed a prompting scheme (see Appendix E) that enabled it to function as an automatic research agent, capable of reading code, generating hypotheses, and proposing modifications grounded in experimental feedback.
@@
We employ GPT-5 (2025-08-07) and Gemini-2.5-Pro (2025-06-17) for TheAIScientist and AIDE. For Claude Code, it is constrained to its native models and therefore we use Opus-4.1 (2025-08-05).
@@ 5.3 Results and Discoveries
Table 2: Comparison of Performance among different agents. G2.5-Pro denotes Gemini-2.5-Pro. $(\uparrow)$ indicates higher is better, $(\downarrow)$ indicates lower is better.
@@
Causality $(\downarrow)$ 1.1445 1.0063 0.9925 0.9683 0.9549 0.9840
Robust. & Rel.()($\uparrow$) 0.4848 0.9311 0.9205 0.9174 0.9633 0.8921
@@
Fair. & Bias $(\downarrow)$ 0.3787 0.0603 0.1002 0.0385 0.0917 0.3787
Table 3: Comparisons of different agents across Exploration Diversity, Step Success Rate, Step Completion Rate, Cost. M stands for Million, H for Hours, and Min for Minutes. AIDE does not support recording token usage.
@@
Exploration Diversity 28.46±9.97 20.86±4.42 28.60±13.32 18.41±12.68 8.75±6.07
Step Success Rate $(\uparrow)$ 0.83±0.27 0.80±0.26 0.64±0.30 0.66±0.34 0.83±0.32
Step Completion Rate $(\uparrow)$ 1.00±0.00 1.00±0.00 0.54±0.38 0.69±0.24 0.07±0.06
Total Token Cons. (M) $(\downarrow)$ 6.04±2.18 5.45±2.51 N/A N/A 9.06±3.64
Step Token Cons. (M) $(\downarrow)$ 0.06±0.02 0.05±0.03 N/A N/A 3.10±4.34
Total Time Cons. (H) $(\downarrow)$ 9.42±5.25 11.74±8.37 5.46±6.92 9.49±7.24 1.18±1.80
Step Time Cons. (Min) $(\downarrow)$ 11.05±10.71 14.23±14.55 12.55±12.17 10.48±8.29 13.26±11.29
5.3.1 Comparison of Agents with Different Research Strategies
As shown in Tab. 2, the combination of TheAIScientist with Gemini-2.5-Pro achieved the best performance, ranking first in 4 out of 8 tasks. The combination of AIDE with Gemini-2.5-Pro ranked second, securing top results in 2 out of the 8 tasks. These findings suggest that TheAIScientist performs better in discovering novel and effective machine learning methods, compared to AIDE and Claude Code.
As illustrated in Fig. 2, TheAIScientist adopts a research exploration strategy that is broad but shallow, while AIDE maintains both medium breadth and depth. In contrast, Claude Code exhibits a narrow yet deep exploration pattern. For a detailed explanation of the research strategies adopted by each agent, refer to Section 5.1. Considering the research exploration strategy together, the results suggest that a broader research exploration space adopted by TheAIScientist is more effective for discovering promising ideas. This insight offers practical guidance for machine learning research agent that broadly exploring diverse ideas could be more productive than focusing on a single direction.
5.3.2 Analysis of Exploration Diversity
As shown in Tab. 2 and Tab. 3, several key findings emerge from our experimental results. First, TheAIScientist and AIDE demonstrate significantly higher exploration diversity compared to Claude Code, while also exhibiting superior overall performance across the tasks. Second, when both systems employ GPT-5 as the underlying language model, TheAIScientist and AIDE achieve comparable diversity scores (28.46 vs. 28.60, respectively), and both methods attain one best result each. Third, when utilizing Gemini-2.5-Pro, TheAIScientist exhibits higher diversity than AIDE (20.85 vs. 18.41) and achieves a greater number of best results (4 vs. 2). These observations suggest that, with the LLM held fixed, agents with greater exploration diversity tend to achieve better outcome performance. In contrast, when the agent is held fixed, exploration diversity is not a decisive factor; performance also depends on other capabilities of different LLMs (e.g., reasoning and code-writing ability).
Furthermore, TheAIScientist achieves a mean exploration diversity of 24.66, which exceeds both AIDE’s 23.50 and Claude Code’s 8.75. This enhanced diversity can potentially be attributed to TheAIScientist’s more exploration breadth-oriented search strategy, as illustrated in Sec. 5.3.1.
Table 4: Correlation analysis between exploration diversity and outcome. When computing the correlation, the outcome scores are negated for tasks where lower values indicate better performance, ensuring that positive correlations consistently represent improved performance with increased exploration diversity.
Metric Gen. Data Eff. Rep. Learn. Cont. Learn. Causality Robust. Privacy Fair.
p-value 0.1137 0.0121 0.1358 0.0002 0.4668 0.1944 0.6479 0.0517
Correlation 0.4418 0.6287 0.4036 0.8374 -0.2036 -0.3846 -0.1286 0.5292
To further validate the relationship between exploration diversity and performance, we computed the correlation coefficients between exploration diversity and performance. As presented in Tab. 4, our findings reveal positive correlations across five tasks, with p-values indicating significant, borderline significant, or suggestive evidence of this relationship ($p<0.05$ for significant; $p<0.10$ for borderline significant; $p<0.15$ for suggestive evidence; $p\geq 0.15$ for not significant). Notably, for these five tasks, higher exploration diversity was associated with better performance.
@@ 5.3.3 Token and Time Consumption
5.3.4 Case Analysis
Worst Case Analysis
TheAIScientist strongly favors algorithmic modifications (56.2%) with zero implementation errors; AIDE presents balanced distribution but higher implementation issues (31.2%); while Claude Code shows elevated parameter configuration rates (37.5%) matching its algorithmic modification rate. Across LLMs, Gemini-2.5-Pro exhibits extremely strong algorithmic modification tendency (68.8%), GPT-5 primarily focuses on structural modifications (50.0%), and Opus-4.1 maintains balanced distribution between parameter configuration and algorithmic modifications. See Appendix B for detailed results.
Best Case Analysis
TheAIScientist shows a strong algorithmic innovation focus, predominantly targeting core algorithmic logic; AIDE combines strong algorithmic innovation with hyperparameter tuning; while Claude Code uniquely favors architectural modifications including layer additions and normalization adjustments. Across LLMs, GPT-5 exclusively prioritizes algorithmic modifications (100.0%), Gemini-2.5-Pro exhibits the most diversified approach with unique data augmentation efforts (28.6%), and Opus-4.1 balances attention between architectural and algorithmic modifications. See Appendix B for detailed results.
@@ 5.3.5 Additional Observations
We found that AIDE sometimes misinterprets the structure and logic of target codebases. In certain cases, it generated new classes or components that were never integrated into the actual execution, resulting in no functional improvement over the baseline. As shown in Tab. 2, AIDE failed to improve the baseline in tasks related to Generalization and Data Efficiency. This may stem from the fact that AIDE only supports iterative modifications on a single file. However, real-world ML research codebases are often complex and span multiple files, making AIDE insufficient for addressing realistic research tasks.
@@
We encountered the issue of early termination of AIDE and Claude Code. For AIDE, the agent sometimes terminated prematurely due to its commercial version Weco which relies on cloud infrastructure that occasionally failed during execution. For Claude Code, early stopping was often triggered by the model’s internal reasoning, where the LLM would decide not to continue even when further actions are possible.
@@ 6 Conclusion
In this work, we introduce FML-bench, a benchmark designed to evaluate ML research agents in scientific research settings. Unlike existing benchmarks that adopt an engineering-oriented perspective through application-oriented tasks and performance-focused evaluation, FML-bench comprises 8 diverse and fundamental ML research problems constructed upon widely recognized academic repositories. We further propose complementary metrics that characterize agent behavior throughout the research process, notably Exploration Diversity, along with Step Success Rate and Step Completion Rate for assessing agent reliability. Through systematic evaluation of state-of-the-art research agents on FML-bench, we find that agents employing broad exploration strategies exhibit higher exploration diversity and achieve superior performance, and that exploration diversity positively correlates with performance improvements across multiple tasks. We hope FML-bench and our findings provide a foundation for evaluating research agents from a scientific research perspective and offer insights for building more effective research agents.
Impact Statement
This work impacts ML applications by shifting the evaluation of research agents from narrow, answer-matching benchmarks toward realistic, process-oriented scientific problem solving, which may lead to agents that generalize better and provide more reliable assistance in real research workflows. While research agents themselves may raise ethical concerns (e.g., misuse or over-automation of scientific judgment), FML-bench does not introduce inherent ethical risks; rather, it serves as a neutral tool for observing and analyzing agent behavior. Moreover, by grounding tasks in realistic research problems and emphasizing the investigation of agent behavior, FML-bench promotes more meaningful progress and greater transparency in research agent evaluation. In addition, our proposed unified input–output interfaces enable extensibility across repositories, datasets, and baselines, which we believe can help encourage diverse, reproducible, and behavior-focused evaluation.
@@ Acknowledgements
Appendix A Task Design Details
Generalization.
@@ Appendix A Task Design Details
This task evaluates the ability to develop algorithms that transfer effectively across domains, which is critical for robust real-world deployment. We selected the ColoredMNIST dataset (Arjovsky et al., 2019), which introduces spurious correlations between color and class labels, creating a controlled yet challenging testbed for distribution shift. As the baseline, we adopted Empirical Risk Minimization (ERM) (Vapnik, 1998), the most fundamental training paradigm that provides a clear reference point for measuring generalization improvements. We constructed this task using the DomainBed repository (Gulrajani and Lopez-Paz, 2020), a widely used framework for domain generalization research. Agents are required to train on a source domain and are evaluated on a held-out target domain under distribution shift. The evaluation metric is out-of-domain accuracy, with the objective of maximizing generalization performance while maintaining in-domain accuracy.
Data Efficiency.
@@
This task measures the ability to enhance few-shot learning performance by optimizing decision rules under tight data constraints. We selected the Mini-ImageNet dataset (Vinyals et al., 2016), a standard few-shot learning benchmark whose fine-grained inter-class similarities and limited training examples per class present meaningful challenges for metric-based classification. As the baseline, we adopted Prototypical Networks (Snell et al., 2017), a classic and widely recognized metric-based approach that computes class prototypes in the embedding space. We constructed this task using the Easy-Few-Shot-Learning repository (Sicara, 2024). Agents must operate under a frozen backbone and propose improved algorithms for metric-based classification in the embedding space. The evaluation metric is few-shot classification accuracy, with the objective of improving the classifier’s ability to generalize to novel classes.
Representation Learning.
This task tests the ability to learn generalizable and semantically meaningful features from unlabeled data through self-supervised learning. We selected the CIFAR-10 dataset (Krizhevsky et al., 2009), which provides a computationally efficient yet sufficiently challenging benchmark for evaluating learned representations. As the baseline, we adopted MoCo (Momentum Contrast) (He et al., 2020), a widely recognized and effective contrastive learning method that maintains a momentum-updated encoder for consistent representations. We constructed this task using the Lightly repository (Lightly-AI, 2025), a modular framework for self-supervised learning. Agents pretrain encoders on the CIFAR-10 training split without labels using a ResNet-18 backbone, and evaluation is conducted via linear probing accuracy with the encoder frozen. The objective is to improve representation quality such that a linear classifier trained on frozen features achieves higher Top-1 accuracy on the test split, while keeping parameter count and FLOPs comparable to the baseline.
Continual Learning.
@@
This task evaluates long-term adaptability in non-stationary environments and the ability to mitigate catastrophic forgetting when learning sequentially across multiple tasks. We selected the splitMNIST dataset (Deng, 2012) under the class-incremental learning scenario, where the model must learn five sequential tasks (each introducing two new digit classes) using a shared ten-way classification head—a setting that rigorously tests knowledge retention. As the baseline, we adopted Synaptic Intelligence (SI) (Zenke et al., 2017), an established regularization-based method that estimates parameter importance online to protect previously learned knowledge. We constructed this task using the Continual-Learning repository (van de Ven et al., 2022). The evaluation metric is average accuracy across all five tasks, with the objective of maximizing performance while avoiding unfair advantages in model size or computational budget.
Causality.
This task assesses the capacity to reason about interventions and estimate treatment effects, which is essential for decision-making in high-stakes environments. We selected the IHDP (Infant Health and Development Program) dataset (Hill, 2011), a semi-synthetic benchmark built from a real randomized controlled trial with simulated treatment effects, enabling ground-truth evaluation of causal inference methods. As the baseline, we adopted DragonNet (Shi et al., 2019), an established neural network architecture designed for treatment effect estimation that jointly models outcome prediction and propensity scoring. We constructed this task using the CausalML repository (Chen et al., 2020). The evaluation metric is mean absolute error in estimating individual treatment effects, with the objective of improving the precision of causal effect estimation.
Robustness and Reliability.
This task probes the ability to improve model robustness against adversarial data corruption while preserving performance on clean data. We constructed a poisoned MNIST dataset incorporating diverse backdoor attack methods, including edge-based triggers and various distributed attack patterns, to create a challenging defense scenario. As the baseline, we adopted dp-instahide (Borgnia et al., 2021), a privacy-preserving defense that mixes inputs with public data and applies differential privacy noise to hinder inversion and poisoning attacks. We constructed this task using the Adversarial Robustness Toolbox (ART) repository (Nicolae et al., 2018). Agents are tasked with proposing defenses that reduce the effectiveness of poisoning attacks while maintaining high clean accuracy. The evaluation metric is a defense score defined as the harmonic mean of clean accuracy and resistance accuracy against backdoor attacks.
Privacy.
This task is critical for assessing the ability to enhance privacy protections against membership inference attacks, which attempt to determine whether specific data points were used in model training. We selected the CIFAR-10 dataset (Krizhevsky et al., 2009) as the training corpus, providing a standard image classification setting for privacy evaluation. As the baseline, we adopted Wide-ResNet-28-2 (Zagoruyko and Komodakis, 2016), a widely used architecture that provides a representative target model for privacy auditing. We constructed this task using the PrivacyMeter repository (Murakonda and Shokri, 2020), which implements both standard Membership Inference Attacks (MIA) and Robust MIA (RMIA) that leverages reference models and population data for more rigorous privacy assessment. The evaluation metric is the attack AUC, with the objective of driving it toward 0.5 (random guessing) while maintaining classification accuracy, thereby reducing information leakage about training data membership.
Fairness and Bias.